3.A laser beam is incident on two slits with a separation of 0.200 mm, and a screen is placed 5.00 m from the slits. An interference pattern appears on the screen. If the angle from the center fringe to the first bright fringe to the side is 0.181 degrees, what is the wavelength of the laser light?
1
Expert's answer
2020-06-15T10:27:39-0400

"d\cdot \sin\alpha=n\lambda \to \lambda=\frac{d\cdot \sin\alpha}{n}=\frac{0.2\cdot10^{-3}\cdot \sin0.181°}{1}=6.32\cdot10^{-7}m=632nm"